A solid disk of 5.0 kg is rotating at a constant angular velocity of 9.5 rad/s.

Calculate the angular momentum of the disk.
Answer #1

Solution) m = 5 kg

w = 9.5 rad/s

Angular momentum , L = ?

L = I×w

I is moment of inertia of solid disk

I = (1/2)(m)(r^2)

L = (1/2)(m)(r^2)(w)

L = (1/2)(5)(r^2)(9.5)

L = 23.75(r^2)

Radius 'r ' value is not given